OpenStudy (anonymous):

3.5^(3x+1)=65.4 How do you solve this using logarithms

OpenStudy (anonymous):

try \[\log _{3.5} ( 3.5 ^{3x + 1} ) = \log _{3.5} 65.4\] to reduce it to \[3x + 1 = \log _{3.5} 65.4\]